Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Tuesday, December 22, 1998

The following is a brief summary of the environmental monitoring activities that were completed on December 22, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

12/22/98 (Tuesday)

Segment 1 report for 12/22/98: Testing of Equipment continued at Emidio pump station. Backfill has been completed on the west side of the Stream Crossing #4 bore (EMP 9.03) and is in progress on the east side. Backfill is also under way at the Tecuya Creek/I-5 bore (EMP 9.83) and a concrete valve box has been placed in the trench there. Tie-in continued on the south side of the California Aqueduct bore (EMP 12.28). Installation of casing continued at the Stream Crossing #9 (EMP 15.35) and Tejon Bench/I-5 (EMP 16.49) bores. Cable was pulled from EMP's 11.77 to 10.72. On Grapevine Rd. backfill continued to EMP 15.78. At Grapevine Pump station electrical work continued. On Tejon Bench, shading continued from EMP's 16.87 to 16.57. Excavation occurred to ~ 17.94. A segment of pipe was welded, x-rayed and coated at EMP 17.97. On Lebec Rd. water was being pumped from the trench and a track hoe began to clear vegetation at EMP 21.71. Crews worked to lower the joint of pipe extending up from the bore at the southern end of the Cuddy Creek Tributary bore (EMP 25.47).

Segment 2C (Castaic to Sylmar): Report for 12-22-98. Crews continued a general clean up of several areas of the right of way. These areas included the yard at Castaic, the extra work space at the north end of Castaic Rd and the work space at Weldon Cyn and Calgrove Blvd. In addition, at the Castaic yard crews began removing tanks and storage sheds. Paving crews progressed with patch work along the right of way from Castaic Rd to Magic Mtn Pkwy.

Segment 3C (Sylmar to San Fernando): Report for 12-22-98. At Balboa, hydro-test crews dewatered the segment 2 line then placed pigs and welded the test head for the segment 3 section. In addition, at the aqueduct overcrossing concrete train barriers were constructed to protect the weak points of the casing from a train derailment. Further, crews poured concrete near the now covered channel and began a general clean up of the rail right of way south of the channel. Near Tuxford crews continued clean up and completed backfill of the bore pit while crews compacted backfill at the Truman St. channel overcrossing. Segment 3A (Taylor yard to San Fernando): Crews pulled fiber optic cable from Paxton St to Victory St in Glendale. At the Cornfield yard, hydro test crews welded and x-rayed a test head then installed the muffler unit.

Inspection Summar

In the past 24 hours, 3 inspection event(s) were reported for 3 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page